# A Survey on Graph Neural Networks for Time Series: Forecasting, Classification, Imputation, and Anomaly Detection

### v1

@@ 9 Future Directions
Pre-training, Transfer Learning, and Large Models. Pre-training, transfer learning, and large models are emerging as potent strategies to bolster the performance of GNNs in time series analysis, especially when data is sparse or diverse. These techniques hinge on utilizing learned representations from one or more domains to enhance performance in other related domains [3, 289]. Despite the challenges, recent advancements, such as Panagopoulos et al.’s model-agnostic meta-learning schema for COVID-19 spread prediction in data-limited cities, and Shao et al.’s pre-training enhanced framework for spatial-temporal GNNs, demonstrate the potential of these strategies [290, 99]. The exploration of pre-training strategies and GNN transferability for time series tasks is a burgeoning research area, especially in the current era of generative AI and large models, which showcase the potential for a single, multimodal model to address diverse tasks [291]. However, several challenges remain, including the limited availability of time series data for large-scale pre-training compared to language data for large language models (LLMs) [292], ensuring the generalizability of learned knowledge to prevent negative transfers, and designing effective pre-training strategies that capture complex spatial-temporal dependencies. Addressing these challenges is pivotal for the future development and application of GNN4TS.
@@
Robustness and Interpretability. Robustness of GNNs refers to their ability to maintain stability under the influence of perturbations, particularly those that are deliberately engineered by adversaries [293]. This quality becomes critical when dealing with time series data generated by rapidly evolving systems. Any operational failures within GNNs can potentially precipitate adverse consequences on the integrity of the entire system [294, 45]. For instance, if a GNN fails to adequately handle noise or perturbations in a smart city application, it might disrupt essential traffic management functions. Similarly, in healthcare applications, the inability of a GNN to remain robust amidst disturbances could lead to healthcare providers missing out on critical treatment periods, potentially having serious health implications for patients. While GNNs have demonstrated superior performance across numerous applications, improving their robustness and creating effective failure management strategies remains vital. This not only enhances their reliability but also widens their potential usage across contexts.
@@
The interpretability of GNNs plays an equally pivotal role in facilitating the transparent and accountable use of these sophisticated tools. This attribute sheds light on the opaque decision-making processes of GNNs, allowing users to comprehend the reasoning behind a given output or prediction. Such understanding fosters trust in the system and enables the discovery of latent patterns within the data [295, 293]. For example, in healthcare and financial time series analyses, interpretability may illuminate causal factors[296, 297], facilitating more informed decision-making. As we strive to harness the full potential of GNN4TS, advancing their interpretability is paramount to ensuring their ethical and judicious application in increasingly complex environments.
@@
Uncertainty Quantification. Time series data, by its nature, is dynamic and often fraught with unpredictable fluctuations and noise. In such environments, the ability of a model to account for and quantify uncertainty can greatly enhance its reliability and utility[298, 299]. Uncertainty quantification provides a probabilistic measure of the confidence in the predictions made by the model, aiding in the understanding of the range and likelihood of potential outcomes[300]. This becomes particularly important when GNNs are used for decision-making processes in fields where high stakes are involved, such as financial forecasting, healthcare monitoring[255, 258], or traffic prediction in smart cities[82, 84, 54, 228]. Despite progress, a gap remains in the current GNN models, which largely provide point estimates[82, 84, 61, 154, 49, 46, 74], inadequately addressing the potential uncertainties. This underlines an essential research direction: developing sophisticated uncertainty quantification methods for GNNs to better navigate the complexities of time series data. This endeavor not only enhances the interpretability and reliability of predictions but also fosters the development of advanced models capable of learning from uncertainty. Thus, uncertainty quantification, albeit nascent, represents a promising and pivotal pathway in the ongoing advancement of GNN4TS.
@@
10 Conclusions
This comprehensive survey bridges the knowledge gap in the field of graph neural networks for time series analysis (GNN4TS) by providing a detailed review of recent advancements and offering a unified taxonomy to categorize existing works from task- and methodology-oriented perspectives. As the first of its kind, it covers a wide range of tasks including forecasting, classification, anomaly detection, and imputation, providing a detailed understanding of the state-of-the-art in GNN4TS. We also delve into the intricacies of spatial and temporal dependencies modeling and overall model architecture, offering a fine-grained classification of individual studies. Highlighting the expanding applications of GNN4TS across various sectors, we demonstrate its versatility and potential for future growth. This survey serves as a valuable resource for machine learning practitioners and domain experts interested in the latest advancements in this field. Lastly, we propose potential future research directions, offering insights to guide and inspire future work in GNN4TS.

### v3

@@ Appendix G Future Directions
Pre-training, Transfer Learning, and Large Models. Pre-training, transfer learning, and large models are emerging as potent strategies to bolster the performance of GNNs in time series analysis [61, 354, 200], especially when data is sparse or diverse. These techniques hinge on utilizing learned representations from one or more domains to enhance performance in other related domains [3, 355]. Recent successful examples are Panagopoulos et al.’s [201] model-agnostic meta-learning schema for COVID-19 spread prediction in data-limited cities, and Shao et al.’s [61] pre-training enhanced framework for spatial-temporal GNNs. The exploration of pre-training strategies and GNN transferability for time series tasks is a burgeoning research area, especially in the current era of generative AI and large models, which showcase the potential for a single, multimodal model to address diverse tasks [202]. However, several challenges remain, including the limited availability of time series data for large-scale pre-training compared to language data for large language models (LLMs) [203], to ensure wide coverage and transferability of learned knowledge, and designing effective pre-training strategies that capture complex spatial-temporal dependencies. Addressing these challenges is pivotal for the future development and application of GNN4TS.
@@
Robustness. Robustness of GNNs refers to their ability to handle various forms of data perturbations and distribution shifts, particularly those that are deliberately engineered by adversaries [204]. This quality becomes critical when dealing with time series data generated by rapidly evolving systems. Any operational failures within GNNs can potentially precipitate adverse consequences on the integrity of the entire system [356, 205]. For instance, if a GNN fails to adequately handle noise or data corruption in a smart city application, it might disrupt essential traffic management functions. Similarly, in healthcare applications, the inability of a GNN to remain robust amidst disturbances could lead to healthcare providers missing out on critical treatment periods, potentially having serious health implications for patients. While GNNs have demonstrated superior performance across numerous applications, improving their robustness and creating effective failure management strategies remains vital. This not only enhances their reliability but also widens their potential usage across contexts.
@@
Interpretability. The interpretability of GNNs plays an equally pivotal role in facilitating the transparent and accountable use of these sophisticated tools. This attribute sheds light on the opaque decision-making processes of GNNs, allowing users to comprehend the reasoning behind a given output or prediction. Such understanding fosters trust in the system and enables the discovery of latent patterns within the data [357, 204]. For example, in drug discovery [32, 358] and financial time series analyses [359, 360], interpretability may illuminate causal factors, facilitating more informed decision-making. As we strive to harness the full potential of GNN4TS, advancing their interpretability is paramount to ensuring their ethical and judicious application in increasingly complex environments.
@@
Uncertainty Quantification. Time series data, by its nature, is often fraught with unpredictable noise and uncertainty associated with the data-generating process. The ability of a model to account for and quantify uncertainty can greatly enhance its reliability and utility[361, 362, 363]. Uncertainty quantification provides a probabilistic measure of the confidence to the predictions made by the model and to the system state estimates [212, 223], aiding in the understanding of the range and likelihood of potential outcomes[364]. This becomes particularly important when GNNs are used for decision-making processes in fields where high stakes are involved, such as financial forecasting, healthcare monitoring[288, 291], or traffic prediction in smart cities[76, 86, 41, 265]. Despite progress, a gap remains in the current GNN models, which largely provide point estimates[76, 86, 51, 89, 36, 33, 63], inadequately addressing the potential uncertainties. This underlines an essential research direction: developing sophisticated uncertainty quantification methods for GNNs to better navigate the complexities of time series data. This endeavor not only enhances the interpretability and reliability of predictions but also fosters the development of advanced models capable of learning from uncertainty. Thus, uncertainty quantification represents a pivotal element in the ongoing advancement of GNN4TS.